The length of a metal wire is drawn and doubled. How many times will the new resistance of the wire become?
1
Expert's answer
2020-12-29T15:32:40-0500

We know that resistance of wire depend on its length and cross sectional area.Mathematically


"R=\rho\frac{l}{A}"

Where R is resistance of wire, l is length of wire ,A is cross sectional area and "\rho" is resistivity of metal.

Note that after stretching wire its cross sectional area and length varies but Its volume"(Al=constant)"remain constant.It means that If length is doubled , automatically its cross sectional area is halved. Thus New resistance of wire


"R_{new}=\rho\frac{{2l}}{\frac{A}{2}}"


"R_{new}=4\rho\frac{l}{A}=4R"

Thus New resistance of wire is 4 times of the initial value.